# Inferring Foresightedness in Dynamic Noncooperative Games

### v2

@@ Abstract
Dynamic game theory is an increasingly popular tool for modeling multi-agent, e.g. human-robot, interactions. Game-theoretic models presume that each agent wishes to minimize a private cost function that depends on others’ actions. These games typically evolve over a fixed time horizon, specifying how far into the future each agent plans. In practical settings, however, decision-makers may vary in foresightedness. We conjecture that quantifying and estimating each agent’s foresightedness from online data will enable safer and more efficient interactions with other agents. To this end, we frame this inference problem as an inverse dynamic game. We consider a specific parametrization of each agent’s objective function that smoothly interpolates myopic and farsighted planning. Games of this form are readily transformed into parametric mixed complementarity problems; we exploit the directional differentiability of solutions to these problems with respect to their hidden parameters to solve for agents’ foresightedness. We conduct two types of experiments: one with synthetically generated pedestrian motion at a crosswalk and the other with real-world intersection data involving people walking, biking, and driving vehicles. The results of these experiments demonstrate that explicitly inferring agents’ foresightedness enables game-theoretic models to more accurately model agents’ behavior. Specifically, our results show a $\mathbf{33\%}$ more accurate prediction of foresighted behavior on average compared to the baseline in real-world scenarios.
@@ I Introduction
Robot planning problems often involve strategic interactions between human and robotic decision makers. Dynamic game theory captures many of these problems’ complexities, such as conflicting goals, coupled strategies, and long-term consequences. However, for an agent to use game theoretic planning algorithms, it must understand other agents’ hidden objectives, or objectives that are not known a priori. To remedy this, we can cast inferring hidden objectives as an inverse game [hadfield2016cooperative, kuleshov2015inverse, molloy2022inverse]. In practice, agents can use algorithms to simultaneously infer others’ objectives and plan trajectories that seamlessly interact with one another over time.
@@
Existing inverse game solvers often assume that decision makers care equally about each moment in time. In practice, both humans and robots can violate this assumption. For example, in an urban driving game modeling a busy intersection, a farsighted driver may slow down further in advance of the intersection than a short-sighted, or myopic, driver would. Robots incorrectly predicting the myopic driver’s foresight may take prematurely evasive actions, resulting in unexpected and unnecessarily risky behavior. Thus, to interact safely and efficiently, autonomous robots must be able to infer the foresightedness of other decision makers.
@@
Figure 1: Representative example showcasing how the proposed formulation improves our ability to model agents’ behavior in noncooperative interactions. Here, two simulated pedestrians are crossing an intersection, and by correctly inferring their degree of foresightedness (i.e., the $\gamma$ parameter in the inset), our method recovers their trajectories than a baseline approach, which does not account for agents’ foresighted decision-making.
@@
This paper aims to address the gap between the assumptions of existing game theoretic algorithms and decision makers’ foresightedness. In particular, existing methods typically model each agent’s objective as a sum of cost functions which can depend upon all agents’ states and actions. In this work, we consider a time-discounted cost formulation, which smoothly interpolates foresight for each agent. Additionally, we develop a gradient-based algorithm for inverting the corresponding game and identifying each agent’s degree of foresightedness. Our specific contributions are:
  1. 1.
    an explicitly foresighted game formulation,
  2. 2.
    an efficient method to infer each agent’s degree of foresightedness from data collected online,
  3. 3.
    and a series of simulated and real-world experiments that demonstrate the benefits of modeling agents’ foresight.
@@
In particular, we perform experiments simulating the interaction of two myopic human agents moving through a crosswalk with intersecting trajectories as shown in Figure 1. We also conduct experiments with real-world data of interactions between humans walking, biking, and driving vehicles, shown in Figure 2. Results demonstrate that our method recovers more accurate models of foresighted agents than existing models, as well as being more robust to noise in partially observable settings. Specifically, our method reduces trajectory error by over $33\%$ on average in comparison to cutting-edge baseline methods.
@@ II-A Noncooperative Dynamic Games
An $N$-agent dynamic game models each agent’s choice of strategy in terms of a optimization problem, whose objective and/or constraints can depend upon other agents’ strategies. Dynamic games admit different solutions depending on the information available to each agent at every time step [BasarOlsder]. In this work, we consider the Nash equilibrium concept in the setting of open-loop strategies, i.e., we presume that at each time $t$ agents only have information about the initial state of the game. Early works in this setting focused primarily on problems with convex, quadratic costs and linear game dynamics, without any additional constraints [starr1969nonzero, BasarOlsder, engwerda1998open, engwerda1998computational]. However, finding a Nash equilibrium in more general settings is often intractable [daskalakis2009complexity]. Consequently, many recent works develop iterative trajectory optimization methods to find approximate, local open-loop Nash equilibria in games with nonlinear costs and complex constraint structures. Such approximations satisfy first-order optimality conditions of the underlying optimization problem for each agent, and second-order conditions for local optimality must generally be checked a posteriori. One class of such methods employs root-finding techniques for the underlying joint $N$-agent Karush-Kuhn-Tucker (KKT) system [cleac2019algames, di2020first, zhu2023sequential]. Another class employs iterative best-response algorithms to approximate Nash equilibria [spica2020real, wang2019game].
@@ II-B Inverse Dynamic Games
Inverse dynamic games seek to identify unknown aspects of game models, e.g., agents’ intentions and constraints, from observed interactions. Early approaches minimize the residual of the KKT system to identify possible game parameter values [ROTHFU201714909, awasthi2020inverse, inga2019inverse]. However, this type of approach requires full state-action demonstration to evaluate the residuals, posing a challenge for online estimation tasks, where not all agents’ states and actions are available. Worse, the assumption that the demonstrations satisfy first-order optimality conditions means that such residual minimization approaches can perform poorly in the presence of observation noise [peters2023ijrr]. To mitigate these issues, recent work formulates inverse game problems in terms of maximum likelihood estimation (MLE), and explicitly impose first-order optimality conditions as constraints [peters2023ijrr, liu2023learningplaytrajectorygames, li2023cost, hu2024think]. In particular, our work builds directly upon this line of work, and specifically extends the solution approach of [liu2023learningplaytrajectorygames], which develops a differentiable game solver that enables inverse game problems to be solved via gradient descent on the unknown game parameters.
While MLE approaches to inverse games demonstrate strong performance in various scenarios, they are fundamentally limited to provide point estimates of unknown parameters. To this end, another line of work seeks to infer full Bayesian posterior distributions for the unknown game parameters, e.g., via particle filtering [peters2020accommodating] and unscented Kalman filtering [le2021lucidgames]. However, high computational costs hinder these approaches from scaling to larger systems. More recently, work in [liu2024autoencodingbayesianinversegames] proposes to embed a differentiable Nash solver into generative models to solve variational inference problems. However, this Bayesian line of inverse game approaches are only approximate in nature, and still suffer from high computational costs.
Finally, we note that inverse games have also been formulated in maximum-entropy settings [mehr2023maxent], inferring opinion dynamics model [hu2024think], and in combination with neural network components [diehl2023energy].
@@ III-A Dynamic Games
A dynamic game [BasarOlsder] is characterized by $N$ agents, with the $i^{\mathrm{th}}$ agent’s control input denoted $u^{i}_{t}\in\mathbb{R}^{m^{i}}\,\,\forall\,\,t$, and a state variable $x_{t}\in\mathbb{R}^{n}$ following given dynamics $x_{t+1}=f_{t}(x_{t},{u}^{1}_{t},{u}^{2}_{t},\dots,{u}^{N}_{t})$ at discrete times $[T]:=\{1,2,...,T\}$. Each agent has a cost function
@@
We shall refer to all agents’ discount factors as a vector $\gamma=(\gamma^{1},\dots,\gamma^{N})\in[0,\infty]^{N}$ for convenience. Observe that the function $\Gamma^{i}(t;\gamma^{i})$ rapidly approaches 00 as $t\to T$ when $\gamma^{i}<1$. Thus, the $i^{\mathrm{th}}$ agent’s actions do not depend upon costs incurred past some effective time horizon $\tilde{T}^{i}\in[T]$.
Thus, we refer to the sequence of game states as $\textbf{x}=(x_{1}^{\top},x_{2}^{\top},\dots,x_{T}^{\top})^{\top}$, agent $i$’s control sequence as $\textbf{u}^{i}=(u^{i,\top}_{1},u^{i,\top}_{2},\dots,u^{i,\top}_{T})^{\top}$, the sequence of all agents’ actions as $\textbf{u}=(\textbf{u}^{1,\top},\textbf{u}^{2,\top},\dots,\textbf{u}^{N,\top})^{\top}$, and all agents’ hidden parameters as $\theta=(\theta^{1,\top},\theta^{2,\top},\dots,\theta^{N,\top})^{\top}$. Then, we can write each cost function as $C^{i}_{t}(x_{t},u^{i}_{t},u^{\neg i}_{t};\theta^{i})=C^{i}_{t}(x_{t},u_{t};\theta^{i})$ and thus agent $i$’s overall cost is $J^{i}(\textbf{x},\textbf{u};\gamma^{i},\theta^{i})$ with a slight abuse of notation. In general, we may also assign each agent a set of inequality constraints $I^{i}(\textbf{x},\textbf{u};\theta^{i})\geq 0$ and a set of equality constraints $E^{i}(\textbf{x},\textbf{u};\theta^{i})=0$. Finally, we can define a game as a tuple:
@@
Note that Equation 2 does not include the dynamics $f_{t}(\cdot)$ explicitly; this is because the dynamics generate $(T-1)$ equality constraints of the form $x_{t+1}-f_{t}(x_{t},{u}^{1}_{t},{u}^{2}_{t},\dots,{u}^{N}_{t})=0$ for each of the $N$ agents. For clarity, we assume that $E^{i}(\cdot)$ contains these terms.
@@ III-C Mixed Complementarity Problem (MiCP)
Finding a generalized Nash equilibrium is computationally intractable [papadimitriou2007complexity]; therefore, it is common to relax the condition in Section III-B to hold only within an open neighborhood of the point $(\textbf{x}^{*},\textbf{u}^{*})$ [mazumdar2019finding, mazumdar2020gradient, heusel2017gans, di2020first, zhu2023sequential]. Such points are called local Nash equilibria [ratliff2016characterization]. Practically, such points can be identified by solving agents’ first order necessary conditions, which constitute a Mixed Complementarity Problem (MiCP)[facchinei_pang]. An MiCP is defined by decision variables $r\in\mathbb{R}^{\eta_{r}},z\in\mathbb{R}^{\eta_{z}}$, as well as functions $c(r,z)$ and $h(r,z)$, such that
$c(r,z)=0$ (4a)
$0\leq z\perp h(r,z)\geq 0$ (4b)
MiCPs can often be solved efficiently via off-the-shelf solvers such as PATH [dirkse1995path].
@@ IV Problem Statement
We presume that an observer has obtained noisy sensor measurements of the game state over time, and denote these observations $\textbf{y}=\big{[}y_{1}(x_{1}),y(x_{2}),y_{3}(x_{3}),\ldots,y_{T}(x_{T})\big{]}$. We assume observations are drawn independently from Gaussian models
$y_{t}(x_{t})\sim\mathcal{N}\big{(}h_{t}(x_{t}),\Sigma_{t}\big{)}$ (5)
with known covariances $\Sigma_{t}$, where $h_{t}(x_{t})$ describes the expected output of the sensor at time $t$.
Given measurements y, we seek to identify game parameters $\hat{\theta},\hat{\gamma}$ that maximizes the likelihood to have observed y from a GOLNE of $\mathcal{G}(\hat{\theta},\hat{\gamma})$. Equivalently, one can minimize covariance-weighted deviations from expected measurements, which we denote $\mathcal{P}$.
@@ V Solution Approach
[…]
@@ V-A Equilibrium Constraint as an MiCP
In this subsection, we show how to convert a GOLNE into a MiCP which encodes its first-order necessary conditions. Similar techniques have been used in [mombaur2010human, englert2017inverse, peters2023online, liu2023learningplaytrajectorygames] to solve inverse optimal control and game problems.
@@
Then, when the gradients of the constraints are linearly independent at a candidate solution point (i.e., the linear independence constraint qualification is satisfied [facchinei_pang, Chapter 3.2]), the following Karush-Kuhn-Tucker (KKT)conditions must hold for each agent $i$:
$\nabla_{\textbf{x}}\mathcal{L}^{i}(\textbf{x},\textbf{u},\lambda^{i},\mu^{i};\theta^{i},\gamma^{i})$ $=0$ (8a)
$\nabla_{u^{i}}\mathcal{L}^{i}(\textbf{x},\textbf{u},\lambda^{i},\mu^{i};\theta^{i},\gamma^{i})$ $=0$ (8b)
$E^{i}(\textbf{x},\textbf{u};\theta^{i})$ $=0$ (8c)
$0\leq\lambda^{i}\perp I^{i}(\textbf{x},\textbf{u};\theta^{i})$ $\geq 0.$ (8d)
We can then structure agents’ joint KKT conditions as a parametric MiCPs Equation 4, in which the primal and dual variables are concatenated as follows:
$r$ $=[\textbf{x}^{\top},\textbf{u}^{\top},\mu^{1,\top},\mu^{2,\top},\dots,\mu^{N,\top}]^{\top}$ (9a)
$z$ $=[\lambda^{1,\top},\lambda^{2,\top},\dots,\lambda^{N,\top}]^{\top}.$ (9b)
For brevity, define $v=(r^{\top},z^{\top})^{\top}$. Then, with a slight abuse of notation, the parameterized MiCP for each agent can be written as described in Equation 4, where functions $c(\cdot)$ and $h(\cdot)$ are defined as
$c(v;\theta,\gamma)=$ $&\begin{bmatrix}\left(\nabla_{\textbf{x}}L^{i}\right)_{i\in[N]}\\ \left(\nabla_{u^{i}}L^{i}\right)_{i\in[N]}\\ \left(E^{i}\right)_{i\in[N]}\end{bmatrix}\end{aligned}$ (10a)
$h(v;\theta,\gamma)=$ $\begin{bmatrix}\bigl{(}I^{i}(v;\theta^{i})\bigr{)}_{i\in[N]}\end{bmatrix}$ (10b)
For brevity, we will define $F(v;\theta,\gamma)=[c(\cdot)^{\top},h(\cdot)^{\top}]^{\top}$. For additional details on connections between GOLNE s and MiCPs in the context of open loop dynamic games, we direct the reader to [facchinei_pang, liu2023learningplaytrajectorygames].

### v3

@@ Abstract
Dynamic game theory is an increasingly popular tool for modeling multi-agent, e.g. human-robot, interactions. Game-theoretic models presume that each agent wishes to minimize a private cost function that depends on others’ actions. These games typically evolve over a fixed time horizon, specifying how far into the future each agent plans. In practical settings, however, decision-makers may vary in foresightedness, or how much they care about their current cost in relation to their past and future costs. We conjecture that quantifying and estimating each agent’s foresightedness from online data will enable safer and more efficient interactions with other agents. To this end, we frame this inference problem as an inverse dynamic game. We consider a specific objective function parametrization that smoothly interpolates myopic and farsighted planning. Games of this form are readily transformed into parametric mixed complementarity problems; we exploit the directional differentiability of solutions to these problems with respect to their hidden parameters to solve for agents’ foresightedness. We conduct three experiments: one with synthetically generated delivery robot motion, one with real-world data involving people walking, biking, and driving vehicles, and one using high-fidelity simulators. The results of these experiments demonstrate that explicitly inferring agents’ foresightedness enables game-theoretic models to make $\mathbf{33\%}$ more accurate models for agents’ behavior.
@@ I Introduction
Robot planning problems often involve strategic interactions between human and robotic decision makers. Dynamic game theory captures many of these problems’ complexities, such as conflicting goals, coupled strategies, and long-term consequences. However, for an agent to use game theoretic planning algorithms, it must understand other agents’ hidden objectives, or objectives that are not known a priori. To remedy this, we can cast inferring hidden objectives as an inverse game [1, 2, 3]. In practice, agents can use algorithms to simultaneously infer others’ objectives and plan trajectories that seamlessly interact with one another over time.
@@
Existing inverse game solvers often assume that decision makers care equally about each moment in time, but in practice, humans and robots weigh future costs differently than present costs. For example, in an urban driving game modeling a busy intersection, a farsighted driver may slow down further in advance of the intersection than a short-sighted, or myopic, driver would. Robots incorrectly predicting the myopic driver’s foresight may take prematurely evasive actions, resulting in unexpected and unnecessarily risky behavior.
@@
Figure 1: Representative example showcasing how the proposed formulation improves our ability to model agents’ behavior in noncooperative interactions. Here, two simulated delivery robots are crossing an intersection, and by correctly inferring their degree of foresightedness (i.e., the $\gamma$ parameter in the inset), our method recovers their trajectories than a baseline approach.
@@
This paper aims to address the gap between the assumptions of existing game theoretic algorithms and decision makers’ foresightedness. In particular, existing methods typically model each agent’s objective as a sum of cost functions which can depend upon all agents’ states and actions. In this work, we consider a time-discounted cost formulation, which smoothly interpolates foresight for each agent. Additionally, we develop a gradient-based algorithm for inverting the corresponding game and identifying each agent’s degree of foresightedness. Our specific contributions are: 1) an explicitly foresighted game formulation, 2) an efficient method to infer each agent’s degree of foresightedness from data collected online, and 3) and a series of simulated and real-world experiments that demonstrate the benefits of modeling agents’ foresight.
@@
In particular, we conduct one simplified experiment with two myopic delivery robots (presented in Figure 1) and two simulated experiments with real-world data (presented in Figure 2 and Figure 3). Results demonstrate that our method recovers more accurate models of foresighted agents than existing models, as well as being more robust to noise in partially observable settings. Specifically, our method reduces trajectory error by over $33\%$ on average in comparison to cutting-edge baseline methods.
@@ II-A Noncooperative Dynamic Games
An $N$-agent dynamic game models each agent’s choice of strategy in terms of a optimization problem, whose objective and/or constraints can depend upon other agents’ strategies. Dynamic games admit different solutions depending on the information available to each agent at every time step [4]. In this work, we consider the Nash equilibrium concept in the setting of open-loop strategies, i.e., we presume that at each time $t$ agents only have information about the initial state of the game. Early works in this setting focused primarily on problems with convex, quadratic costs and linear game dynamics, without any additional constraints [5, 4, 6, 7]. However, finding a Nash equilibrium in more general settings is often intractable [8]. Consequently, many recent works develop iterative trajectory optimization methods to find approximate, local open-loop Nash equilibria in games with nonlinear costs and complex constraint structures. Such approximations satisfy first-order optimality conditions of the underlying optimization problem for each agent, and second-order conditions for local optimality must generally be checked a posteriori. Some well-known methods include root-finding techniques for the underlying joint $N$-agent Karush-Kuhn-Tucker (KKT) system [9, 10, 11] and iterative best-response algorithms to identify local Nash equilibria [12, 13].
@@ II-B Inverse Dynamic Games
Inverse dynamic games seek to identify unknown aspects of game models, e.g., agents’ intentions and constraints, from observed interactions. Early approaches minimize the residual of the KKT system to identify possible game parameter values [14, 15, 16]. However, this type of approach requires full state-action demonstration to evaluate the residuals, posing a challenge for online estimation tasks, where not all agents’ states and actions are available. Worse, the assumption that the demonstrations satisfy first-order optimality conditions means that such residual minimization approaches can perform poorly in the presence of observation noise [17]. To mitigate these issues, recent work formulates inverse game problems in terms of maximum likelihood estimation (MLE), and explicitly impose first-order optimality conditions as constraints [17, 18, 19, 20]. In particular, our work builds directly upon this line of work, and specifically extends the solution approach of [18], which develops a differentiable game solver that enables inverse game problems to be solved via gradient descent on the unknown game parameters.
While MLE approaches to inverse games demonstrate strong performance in various scenarios, they are fundamentally limited to providing point estimates of unknown parameters. To this end, another line of work seeks to infer full Bayesian posterior distributions for the unknown game parameters, e.g., via particle filtering [21] and unscented Kalman filtering [22]. More recently, work in [23] proposes to embed a differentiable Nash solver into generative models to solve variational inference problems. However, high computational costs hinder these approaches from scaling to larger systems.
Finally, we note that inverse games have also been formulated in maximum-entropy settings [24], inferring opinion dynamics model [20], and in combination with neural network components [25].
II-C Non-Game Theoretic Multi-Agent Modeling
Many non-game theoretic approaches have been proposed for modeling multi-agent robotic systems. Social forces have been used to model multi-agent interactions, where agent objectives are introduced through virtual forces (e.g., spring forces to attract an agent) [26, 27]. Potential fields have been used similarly; instead of virtual forces to encode goals and obstacles, the method uses potential fields to repel and attract agents [28]. Relevant works model multi-agent behavior by encoding agent objectives in velocity space rather than position space with velocity obstacles [29, 30]. Additionally, recent work applies machine learning to predict human driving and pedestrian behavior. Maximum entropy inverse reinforcement learning [31, 32] and history-dependent LSTM-based models [33, 34] are examples of learning methods that can forecast vehicle and pedestrian trajectories. Finally, recent works use generative trajectory models to predict pedestrian behavior [35, 36]. Unlike game-theoretic methods, these methods do not explicitly reason about coupling effects in agents’ decision-making. Additionally, it should be noted that our approach exhibits the interpretability of classical methods, such as social force models, but can flexibly accommodate arbitrary specifications of agents’ objective functions, similar to inverse reinforcement learning methods.
@@ III-A Dynamic Games
A dynamic game [4] is characterized by $N$ agents, with the $i^{\mathrm{th}}$ agent’s control input denoted $u^{i}_{t}\in\mathbb{R}^{m^{i}}$ for all discrete times $t\in[T]:=\{1,2,...,T\}$ and a joint state variable $x_{t}\in\mathbb{R}^{n}$ following given dynamics $x_{t+1}=f_{t}(x_{t},{u}^{1}_{t},{u}^{2}_{t},\dots,{u}^{N}_{t})$. Each agent has a cost function
@@
We shall refer to all agents’ discount factors as a vector $\gamma=(\gamma^{1},\dots,\gamma^{N})\in[0,\infty]^{N}$ for convenience. Observe that the function $\Gamma^{i}(t;\gamma^{i})$ rapidly approaches 0 as $t\to T$ when $\gamma^{i}<1$. Thus, the $i^{\mathrm{th}}$ agent’s actions do not depend upon costs incurred past some effective time horizon $\tilde{T}^{i}\in[T]$.
[…]
@@
Note that Equation 2 does not include the dynamics $f_{t}(\cdot)$ explicitly; this is because the dynamics generate (T1)(T-1) equality constraints of the form $x_{t+1}-f_{t}(x_{t},{u}^{1}_{t},{u}^{2}_{t},\dots,{u}^{N}_{t})=0$ for each of the $N$ agents. For clarity, we assume that $E^{i}(\cdot)$ contains these terms.
@@ III-C Mixed Complementarity Problem (MiCP)
Finding a generalized Nash equilibrium is computationally intractable [37]; therefore, it is common to relax the condition in Section III-B to hold only within an open neighborhood of the point $(\textbf{x}^{*},\textbf{u}^{*})$ [38, 39, 40, 10, 11]. Such points are called local Nash equilibria [41]. Practically, such points can be identified by solving agents’ first order necessary conditions, which constitute a Mixed Complementarity Problem (MiCP) [42]. An MiCP is defined by decision variables $r\in\mathbb{R}^{\eta_{r}},z\in\mathbb{R}^{\eta_{z}}$, as well as functions $c(r,z)$ and $h(r,z)$, such that $c(r,z)=0\text{ and }0\leq z\perp h(r,z)\geq 0.$ MiCPs can often be solved efficiently via off-the-shelf solvers such as PATH [43].
@@ IV Problem Statement
We presume that an observer has obtained noisy sensor measurements of the game state over time, and denote these observations $\textbf{y}=\big[y_{1}(x_{1}),y(x_{2}),y_{3}(x_{3}),\ldots,y_{T}(x_{T})\big]$. We assume observations are drawn independently from Gaussian models $y_{t}(x_{t})\sim\mathcal{N}\big(h_{t}(x_{t}),\Sigma_{t}\big)$ with known covariances $\Sigma_{t}$, where $h_{t}(x_{t})$ describes the expected output of the sensor at time $t$.
[…]
@@ V Solution Approach
In this section, we present a constrained gradient descent algorithm for identifying unknown parameters $\theta,\gamma$ in 1. This technique will require us to take derivatives of solutions $(\textbf{x},\textbf{u})$ to Equation 4b with respect to parameters $(\theta,\gamma)$. Therefore, we begin with a discussion of transforming a GOLNE into an MiCP, whose solutions are directionally differentiable with respect to problem parameters [42, Ch. 5].
@@ V-A Equilibrium Constraint as an MiCP
In this subsection, we show how to convert a GOLNE into a MiCP which encodes its first-order necessary conditions. Similar techniques have been used in [44, 45, 46, 18] to solve inverse optimal control and game problems.
@@
Then, when the gradients of the constraints are linearly independent at a candidate solution point (i.e., the linear independence constraint qualification is satisfied [42, Chapter 3.2]), the following Karush-Kuhn-Tucker (KKT) conditions must hold for each agent $i$:
$\nabla_{\textbf{x}}\mathcal{L}^{i}=0,\nabla_{u^{i}}\mathcal{L}^{i}=0,E^{i}$ $=0$ (6a)
$0\leq\lambda^{i}\perp I^{i}(\textbf{x},\textbf{u};\theta^{i})$ $\geq 0.$ (6b)
We can then structure agents’ joint KKT conditions as a parametric MiCP, in which the primal and dual variables are concatenated as $r=(\textbf{x}^{\top},\textbf{u}^{\top},\mu^{1,\top},\mu^{2,\top},\dots,\mu^{N,\top})^{\top}$ and $z=(\lambda^{1,\top},\lambda^{2,\top},\dots,\lambda^{N,\top})^{\top}$. For brevity, define $v=(r^{\top},z^{\top})^{\top}$. Then, with a slight abuse of notation, the parameterized MiCP for each agent can be written as described in Section III-C, where $c(v;\theta,\gamma)=[\left(\nabla_{\textbf{x}}L^{i}\right)^{\top}_{i\in[N]},\left(\nabla_{\textbf{u}^{i}}L^{i}\right)^{\top}_{i\in[N]},\left(E^{i}\right)^{\top}_{i\in[N]}]^{\top}$ and $h(v;\theta,\gamma)=[\bigl(I^{i}\bigr)_{i\in[N]}]^{\top}$. For brevity, we will define $F(v;\theta,\gamma)=[c(\cdot)^{\top},h(\cdot)^{\top}]^{\top}$. For additional details on connections between GOLNE s and MiCPs in the context of open loop dynamic games, we direct the reader to [42, 18].